### v1

@@ 2 Proofs
Let $uv\in E(J)$. Note that, for every perfect matching $M_{J}$ of $J$ that contains $uv$, there is a corresponding perfect matching $M_{H}$ of $H$ that contains all edges of $P_{uv}$ of odd parity. Similarly, for each perfect matching $M_{J}$ of $J$ that does not contain $uv$, there is a corresponding perfect matching $M_{H}$ of $H$ that contains all edges of $P_{uv}$ of even parity. In fact, every perfect matching of $H$ arises in this manner, and $M_{J}\rightarrow M_{H}$ is a bijection between the sets of perfect matchings of $J$ and those of $H$.
Now, suppose that $H$ is a conformal subgraph of a matching covered graph $G$, and let $C$ denote any cut of $G$. For an edge $uv$ of $J$, we say that the corresponding path $P_{uv}$ of $H$ is a $C$-crossing path if at least one edge of $P_{uv}$ lies in $C$, and we call such an edge a $C$-crossing edge. If there is one such edge, we say that $P_{uv}$ crosses $C$ once; if there are two such edges, we say that $P_{uv}$ crosses $C$ twice; and so on. We now state some observations due to Kothari and Murty [9]. They imposed stronger conditions on $J$; however, their proof goes through exactly as is with the weaker assumptions stated below.
@@
In all cases discussed above, observe that the subgraph $H^{\prime}:=G^{\prime}\cap E(H)$ is a bisubdivision of $J$ in $G^{\prime}$, and $M^{\prime}:=M\cap E(G^{\prime})$ is a perfect matching of $G^{\prime}-V(H^{\prime})$. This completes the proof of Corollary 2.2. ∎
@@
Let $H$ denote a bisubdivision of a matching covered graph $J$, where $\delta(J)\geq 3$, that is conformal in a matching covered graph $G$, and let $B$ denote a barrier of $G$. If $G-B$ has distinct components, say $K$ and $L$, each of which contains exactly one branch vertex of $H$, say $u$ and $v$, respectively, then $u$ and $v$ are nonadjacent in $J$.
@@
Note that if $P_{uv}$ meets a component $F$ of $G-B$ that is distinct from $K$ and $L$, then by Proposition 2.1 (i), $P_{uv}$ has precisely two $\partial(F)$-crossing edges and they are of opposite parity; thus, the subpath $P_{uv}\cap F$ is even. Using this observation, and the fact that $B$ is a stable set, we conclude that the subpath $e_{1}P_{uv}e_{2}$ is an even path; this contradicts what we have already established above. ∎
@@
Let $B$ denote a barrier of a matching covered graph $G$, and let $z$ denote an isolated vertex of $G-B$. Then, any three edges in $\partial(z)$ participate in a conformal bisubdivision of $\theta$.
@@
We proceed by induction on the order of $G$. Let $K$ denote a set of three edges incident at $z$. If each (odd) component of $G-B$ is trivial, then $G$ is bipartite, and using Proposition 1.10 we observe that $K$ participates in a conformal bisubdivision of $\theta$.
Now let $L$ be a nontrivial odd component of $G-B$. Let $U:=V(L)$, let $L_{1}:=G/U\rightarrow u$, and let $L_{2}:=G/\overline{U}\rightarrow\overline{u}$. Observe that, in $L_{1}$, the set $B$ is a barrier, and $z$ is an isolated vertex of $L_{1}-B$. By the induction hypothesis, $K$ participates in a conformal bisubdivision of $\theta$, say $H_{1}$, in $L_{1}$; let $M_{1}$ be a perfect matching of $L_{1}-V(H_{1})$. By Lemma 2.3, the branch vertex of $H_{1}$, distinct from $z$, lies in $B$; in particular, $u$ is not a branch vertex of $H_{1}$.
Now, let $F:=E(H_{1})\cap\partial(U)$. It follows from the above, and Proposition 2.1 (i) and (ii), that $|F|\in\{0,2\}$; furthermore, if $|F|=2$, then the two edges belong to the same $\partial(U)$-crossing path of $H_{1}$, and are of opposite parities. If $|F|=0$, let $H_{2}$ denote the null subgraph of $L_{2}$ and let $M_{2}$ denote a perfect matching of $L_{2}$ that contains the unique edge in $M_{1}\cap\partial(U)$. On the other hand, if $|F|=2$, we invoke Proposition 1.5, and let $H_{2}$ denote a conformal cycle in $L_{2}$ containing $F$, and let $M_{2}$ denote a perfect matching of $L_{2}-V(H_{2})$. Observe that, in each case, $H:=H_{1}\cup H_{2}$ is a bisubdivision of $\theta$ in $G$, and $M:=M_{1}\cup M_{2}$ is a perfect matching of $G-V(H)$. ∎
@@
Let $B$ denote a barrier of a matching covered graph $G$, and let $J$ be a matching covered graph with $\Delta(J)\leq 3$. If there exists an (odd) component $L$ of $G-B$ such that $G/\overline{V(L)}$ is $J$-based, then $G$ is also $J$-based.
@@
Depending on the cardinality of $F$, we will define a conformal subgraph $H_{2}$ in $L_{2}$, and a perfect matching $M_{2}$ of $L_{2}-V(H_{2})$. If $F=\emptyset$, we let $H_{2}$ be the null subgraph of $L_{2}$, and we let $M_{2}$ denote a perfect matching of $L_{2}$ that contains the unique edge in $M_{1}\cap\partial(Z)$. If $|F|=2$, we invoke Proposition 1.5, and let $H_{2}$ denote a conformal cycle in $L_{2}$ containing $F$, and let $M_{2}$ be any perfect matching of $L_{2}-V(H_{2})$. If $|F|=3$, we invoke Proposition 2.4, and let $H_{2}$ denote a conformal bisubdivision of $\theta$ containing $F$ in $L_{2}$, and let $M_{2}$ be any perfect matching of $L_{2}-V(H_{2})$. In all three cases, the reader may verify that $H:=H_{1}\cup H_{2}$ is a bisubdivision of $J$ in $G$, and that $M:=M_{1}\cup M_{2}$ is a perfect matching of $G-V(H)$. ∎
@@
If $J$ is any $2$-connected cubic graph, then $J-e_{1}-e_{2}$ is matchable for each pair of edges $e_{1}$ and $e_{2}$.
@@
In a $3$-connected bicritical graph $J$, for any nontrivial cut $C$ and edge $e\in C$, there exists a perfect matching $M_{J}$ such that $|M_{J}\cap(C-e)|\geq 2$.
We now prove a consequence of the above two lemmas. Recall from Section 1.2.2 that every bicritical cubic graph is $3$-connected.
@@
We begin by making an observation that is applicable when $Y$ and $\overline{Y}$ are both nonempty. Since $J$ is $3$-connected as noted earlier, $H$ has at least three $C$-crossing paths; by Proposition 2.1 (i) and (ii), each of them crosses $C$ in exactly one edge. If two of these $C$-crossing edges, say $e_{1}$ and $e_{2}$, are of even parity, then Lemma 2.6 implies that there exists a perfect matching $M_{H}$ of $H$ that contains both $e_{1}$ and $e_{2}$, and thus $|M_{H}\cap C|\geq 2$, contrary to our first observation. Hence, we conclude that at most one $C$-crossing edge is of even parity. If there is such an edge $e$, we let $P_{uv}$ denote the $C$-crossing path that contains $e$ where $uv\in E(J)\cap\partial_{J}(Y)$. Otherwise, we let $uv$ be any edge of $E(J)\cap\partial_{J}(Y)$.
Suppose to the contrary that $|Y|\geq 2$ and $|\overline{Y}|\geq 2$. By Lemma 2.7, there exists a perfect matching $M_{J}$ of $J$ satisfying $|M_{J}\cap(\partial_{J}(Y)-uv)|\geq 2$. Let $M_{H}$ denote the corresponding perfect matching of $H$. By our choice of $uv$, we infer that $|M_{H}\cap C|\geq 2$, contrary to our first observation. This completes the proof of Corollary 2.8. ∎
@@
Let $J$ denote a bicritical cubic graph, and let $B$ denote a barrier of a matching covered graph $G$. If $G$ is $J$-based then $G/\overline{V(L)}$ is $J$-based for some component $L$ of $G-B$.
@@
Since $J-u-v$ has a perfect matching for each pair $u,v\in V(J)$, the reader may observe that $H-u-v$ has a perfect matching, and by conformality of $H$, the graph $G-u-v$ has a perfect matching. Consequently, by Proposition 1.2, at most one of $u$ and $v$ lies in $B$. ∎
@@
If there exist two or more components of $G-B$ that contain branch vertices of $H$, then there exists a component of $G-B$ that contains at least two branch vertices of $H$.
@@
Suppose to the contrary that at least two components of $G-B$ contain branch vertices of $H$, but each such component contains exactly one branch vertex. Note that either $J=\theta$ or otherwise each vertex of $J$ has at least two neighbors. Thus, it follows from 2.9.1 that there exist components $K$ and $L$ of $G-B$ that contain branch vertices $u$ and $v$, respectively, such that $u$ and $v$ are adjacent in $J$. This contradicts Lemma 2.3. ∎
We remark that if $J=\theta$, then the above statement implies that either both branch vertices belong to a particular component of $G-B$, or otherwise precisely one of them lies in $B$.
@@
There exists a component $L$ of $G-B$ that contains at least $|V(J)|-1$ branch vertices.
@@
The statement holds trivially for $J=\theta$. Otherwise, using 2.9.2, there exists a component $L$ of $G-B$ that contains at least two branch vertices of $H$. Observe that $\partial(L)$ is a nontrivial tight cut; since $J$ is bicritical and cubic, we invoke Corollary 2.8 to infer that $\overline{V(L)}$ contains at most one branch vertex. ∎
@@
2.2 Handling $2$-separations: a proof of Theorem 1.25
@@
Let $J$ and $G$ be matching covered graphs, and let $S$ denote a $2$-separation of $G$. If some marked $S$-component of $G$ is $J$-based, then $G$ is also $J$-based.
@@
Let $S:=\{u,v\}$, and let $K$ and $L$ denote two distinct (even) components of $G-S$. Suppose the marked $S$-component $K^{\prime}$ (corresponding to $K$) is $J$-based, and let $H^{\prime}$ denote a conformal bisubdivision of $J$ in $K^{\prime}$. Let $e$ denote the marker edge of $K^{\prime}$, and let $P$ denote a conformal odd $uv$-path in $L$ provided by Proposition 1.22 (iii). Depending on whether or not $e$ belongs to $H^{\prime}$, either $H:=H^{\prime}+P-e$ or $H:=H^{\prime}$, respectively, is a conformal bisubdivision of $J$ in $G$. ∎
@@
Let $S:=\{u,v\}$ denote a $2$-separation of a matching covered graph $G$, and let $J$ denote any $3$-connected cubic graph. Assume that if $J=\theta$ then $G-S$ has precisely two components. If $G$ is $J$-based then some marked $S$-component of $G$ is also $J$-based.
@@
We let $H$ denote a bisubdivision of $J$ that is conformal in $G$. Since $H$ is a subdivision of a $3$-connected graph, and since $S$ is a $2$-separation of $G$, we immediately conclude that there do not exist two components of $G-S$ each of which contains a branch vertex of $H$. This proves the following.
@@
Since $\theta$ is the only $3$-connected cubic graph of order two, if $J\neq\theta$ then there exists a unique marked $S$-component of $G$, denoted by $L^{\prime}$, that contains all branch vertices of $H$. On the other hand, if $J=\theta$ then, by our hypothesis, there are precisely two marked $S$-components; in this case, each of them may contain both branch vertices, and this happens if and only if $u$ and $v$ are the branch vertices of $H$. So, if $J=\theta$, we let $L^{\prime}$ denote the marked $S$-component that minimizes the number of $\partial_{G}(V(L^{\prime}))$-crossing paths. We let $e$ denote the marker edge of $L^{\prime}$, let $D$ denote the even cut $\partial_{G}(V(L^{\prime}))$, and let $M$ denote a perfect matching of $G-V(H)$.
@@
Suppose to the contrary that $H$ has at least two $D$-crossing paths, say $P$ and $Q$. Since all branch vertices belong to $V(L^{\prime})$, and since $S:=\{u,v\}$ is a $2$-separation of $G$, each of $P$ and $Q$ contains $u$ as well as $v$. This immediately implies that $u$ and $v$ are branch vertices of $H$. Furthermore, since $J$ is $3$-connected and cubic, we conclude that $J=\theta$. Consequently, $H=P\cup Q\cup R$, where $R$ is the third $uv$-path. Observe that if $K^{\prime}$ is the marked $S$-component of $G$ distinct from $L^{\prime}$, then the number of $\partial_{G}(V(K^{\prime}))$-crossing paths is at most one, and this contradicts our choice of $L^{\prime}$. ∎
@@
An edge $e$ of a matching covered graph $G$ is removable if $G-e$ is also matching covered. In this case, if $G$ is a brick then $G-e$ need not be a brick; in particular, $G-e$ may have one or two vertices of degree two. In order to recover a brick, at the very least, we need to get rid of the degree two vertices (if any).
@@
Now, let $e$ be a removable edge of a brick $G$. As noted earlier, $G-e$ has zero, one or two vertices of degree two. The graph $H$ obtained from $G-e$ by bicontracting each of these degree two vertices is called the retract of $G-e$. Observe that $|V(G)|-|V(H)|\in\{0,2,4\}$; furthermore, each contraction vertex of $H$ (resulting from the bicontraction of degree two vertices), if any, has degree at least four. We say that $e$ is a thin edge of the brick $G$ if the retract of $G-e$ is also a brick. For instance, the bicorn shown in Figure 8 has a unique removable edge that is also thin. On the other hand, if $e$ is any edge of the Petersen graph $\mathbb{P}$, the retract of $\mathbb{P}-e$ is precisely $T_{6}$; consequently, $\mathbb{P}$ is devoid of thin edges.
@@
In their work [4], Carvalho, Lucchesi and Murty defined four expansion operations — each of which comprises a sequence of zero, one or two bisplittings followed by adding an edge. In particular, if $e$ is a thin edge of a brick $G$, and $H$ is the retract of $G-e$, then $G$ may be obtained from $H$ using one of the four expansion operations. In our work, we do not require the precise definitions of these operations; the interested reader may refer to [4]. We shall simply use the facts that the constituent operations of adding an edge as well as of bisplitting preserve the $\theta$-based property; the former is easy to see, whereas the latter follows from Corollary 2.12. We are now ready to prove our characterization of $\theta$-free bricks restated below.
@@
By Proposition 1.12, $\mathbb{P}$ is $\theta$-free. Clearly, $K_{4}$ is $\theta$-free. Conversely, let $G$ be a $\theta$-free brick. We proceed by induction on the size of $G$. If $G\in\{K_{4},\overline{C_{6}},\mathbb{P}\}$, since $\overline{C_{6}}$ is $\theta$-based, the desired conclusion holds. Otherwise, by the Thin Edge Theorem (2.13), $G$ has a thin edge, say $e$, and let $H$ denote the retract of $G-e$. Clearly, $G-e$ is $\theta$-free; by Corollary 2.12, $H$ is also $\theta$-free. Since $H$ is a brick, by the induction hypothesis, $H\in\{K_{4},\mathbb{P}\}$. Furthermore, $G$ is obtained from $H$ by one of the four expansion operations. Since $H$ is cubic, $G$ is obtained from $H$ by adding an edge. If $H$ is $K_{4}$, clearly $G$ is $\theta$-based; if $H$ is $\mathbb{P}$, by Proposition 1.12, $G$ is $\theta$-based; in either case, we have a contradiction. This proves Theorem 1.27. ∎

### v2

@@ 2 Proofs
Let $uv\in E(J)$. Note that, for every perfect matching $M_{J}$ of $J$ that contains uvuv, there is a corresponding perfect matching $M_{H}$ of $H$ that contains all edges of $P_{uv}$ of odd parity. Similarly, for each perfect matching $M_{J}$ of $J$ that does not contain uvuv, there is a corresponding perfect matching $M_{H}$ of $H$ that contains all edges of $P_{uv}$ of even parity. In fact, every perfect matching of $H$ arises in this manner, and $M_{J}\rightarrow M_{H}$ is a bijection between the sets of perfect matchings of $J$ and those of $H$.
Now, suppose that $H$ is a conformal subgraph of a matching covered graph $G$, and let $C$ denote any cut of $G$. For an edge uvuv of $J$, we say that the corresponding path $P_{uv}$ of $H$ is a $C$-crossing path if at least one edge of $P_{uv}$ lies in $C$, and we call such an edge a $C$-crossing edge. If there is one such edge, we say that $P_{uv}$ crosses $C$ once; if there are two such edges, we say that $P_{uv}$ crosses $C$ twice; and so on. We now state some observations due to Kothari and Murty [9]. They imposed stronger conditions on $J$; however, their proof goes through exactly as is with the weaker assumptions stated below.
@@
In all cases discussed above, observe that the subgraph $H^{\prime}:=G^{\prime}\cap E(H)$ is a bisubdivision of $J$ in $G^{\prime}$, and $M^{\prime}:=M\cap E(G^{\prime})$ is a perfect matching of $G^{\prime}-V(H^{\prime})$. This completes the proof of Corollary $2.$ ∎
@@
Let $H$ denote a bisubdivision of a matching covered graph $J$, where $\delta(J)\geq 3$, that is conformal in a matching covered graph $G$, and let $B$ denote a barrier of $G$. If GBG-B has distinct components, say $K$ and $L$, each of which contains exactly one branch vertex of $H$, say $u$ and $v$, respectively, then $u$ and $v$ are nonadjacent in $J$.
@@
Note that if $P_{uv}$ meets a component $F$ of GBG-B that is distinct from $K$ and $L$, then by Proposition 2.1 (i), $P_{uv}$ has precisely two $\partial(F)$-crossing edges and they are of opposite parity; thus, the subpath $P_{uv}\cap F$ is even. Using this observation, and the fact that $B$ is a stable set, we conclude that the subpath $e_{1}P_{uv}e_{2}$ is an even path; this contradicts what we have already established above. ∎
@@
Let $B$ denote a barrier of a matching covered graph $G$, and let $z$ denote an isolated vertex of GBG-B. Then, any three edges in $\partial(z)$ participate in a conformal bisubdivision of $\theta$.
@@
We proceed by induction on the order of $G$. Let $K$ denote a set of three edges incident at $z$. If each (odd) component of GBG-B is trivial, then $G$ is bipartite, and using Proposition 1.10 we observe that $K$ participates in a conformal bisubdivision of $\theta$.
Now let $L$ be a nontrivial odd component of GBG-B. Let $U:=V(L)$, let $L_{1}:=G/U\rightarrow u$, and let $L_{2}:=G/\overline{U}\rightarrow\overline{u}$. Observe that, in $L_{1}$, the set $B$ is a barrier, and $z$ is an isolated vertex of $L_{1}-B$. By the induction hypothesis, $K$ participates in a conformal bisubdivision of $\theta$, say $H_{1}$, in $L_{1}$; let $M_{1}$ be a perfect matching of $L_{1}-V(H_{1})$. By Lemma 2.3, the branch vertex of $H_{1}$, distinct from $z$, lies in $B$; in particular, $u$ is not a branch vertex of $H_{1}$.
Now, let $F:=E(H_{1})\cap\partial(U)$. It follows from the above, and Proposition 2.1 (i) and (ii), that $|F|\in\{0,2\}$; furthermore, if $|F|=2$, then the two edges belong to the same $\partial(U)$-crossing path of $H_{1}$, and are of opposite parities. If $|F|=0$, let $H_{2}$ denote the null subgraph of $L_{2}$ and let $M_{2}$ denote a perfect matching of $L_{2}$ that contains the unique edge in $M_{1}\cap\partial(U)$. On the other hand, if $|F|=2$, we invoke Proposition 1.5, and let $H_{2}$ denote a conformal cycle in $L_{2}$ containing $F$, and let $M_{2}$ denote a perfect matching of $L_{2}-V(H_{2})$. Observe that, in each case, $H:=H_{1}\cup H_{2}$ is a bisubdivision of $\theta$ in $G$, and $M:=M_{1}\cup M_{2}$ is a perfect matching of GV(H)G-V(H). ∎
@@
Let $B$ denote a barrier of a matching covered graph $G$, and let $J$ be a matching covered graph with $\Delta(J)\leq 3$. If there exists an (odd) component $L$ of GBG-B such that $G/\overline{V(L)}$ is $J$-based, then $G$ is also $J$-based.
@@
Depending on the cardinality of $F$, we will define a conformal subgraph $H_{2}$ in $L_{2}$, and a perfect matching $M_{2}$ of $L_{2}-V(H_{2})$. If $F=\emptyset$, we let $H_{2}$ be the null subgraph of $L_{2}$, and we let $M_{2}$ denote a perfect matching of $L_{2}$ that contains the unique edge in $M_{1}\cap\partial(Z)$. If $|F|=2$, we invoke Proposition 1.5, and let $H_{2}$ denote a conformal cycle in $L_{2}$ containing $F$, and let $M_{2}$ be any perfect matching of $L_{2}-V(H_{2})$. If $|F|=3$, we invoke Proposition 2.4, and let $H_{2}$ denote a conformal bisubdivision of $\theta$ containing $F$ in $L_{2}$, and let $M_{2}$ be any perfect matching of $L_{2}-V(H_{2})$. In all three cases, the reader may verify that $H:=H_{1}\cup H_{2}$ is a bisubdivision of $J$ in $G$, and that $M:=M_{1}\cup M_{2}$ is a perfect matching of GV(H)G-V(H). ∎
@@
If $J$ is any 22-connected cubic graph, then $J-e_{1}-e_{2}$ is matchable for each pair of edges $e_{1}$ and $e_{2}$.
@@
In a 33-connected bicritical graph $J$, for any nontrivial cut $C$ and edge $e\in C$, there exists a perfect matching $M_{J}$ such that $|M_{J}\cap(C-e)|\geq 2$.
We now prove a consequence of the above two lemmas. Recall from Section 1.2.2 that every bicritical cubic graph is 33-connected.
@@
We begin by making an observation that is applicable when $Y$ and $\overline{Y}$ are both nonempty. Since $J$ is 33-connected as noted earlier, $H$ has at least three $C$-crossing paths; by Proposition 2.1 (i) and (ii), each of them crosses $C$ in exactly one edge. If two of these $C$-crossing edges, say $e_{1}$ and $e_{2}$, are of even parity, then Lemma 2.6 implies that there exists a perfect matching $M_{H}$ of $H$ that contains both $e_{1}$ and $e_{2}$, and thus $|M_{H}\cap C|\geq 2$, contrary to our first observation. Hence, we conclude that at most one $C$-crossing edge is of even parity. If there is such an edge $e$, we let $P_{uv}$ denote the $C$-crossing path that contains $e$ where $uv\in E(J)\cap\partial_{J}(Y)$. Otherwise, we let uvuv be any edge of $E(J)\cap\partial_{J}(Y)$.
Suppose to the contrary that $|Y|\geq 2$ and $|\overline{Y}|\geq 2$. By Lemma 2.7, there exists a perfect matching $M_{J}$ of $J$ satisfying $|M_{J}\cap(\partial_{J}(Y)-uv)|\geq 2$. Let $M_{H}$ denote the corresponding perfect matching of $H$. By our choice of uvuv, we infer that $|M_{H}\cap C|\geq 2$, contrary to our first observation. This completes the proof of Corollary 2.8. ∎
@@
Let $J$ denote a bicritical cubic graph, and let $B$ denote a barrier of a matching covered graph $G$. If $G$ is $J$-based then $G/\overline{V(L)}$ is $J$-based for some component $L$ of GBG-B.
@@
Since JuvJ-u-v has a perfect matching for each pair $u,v\in V(J)$, the reader may observe that HuvH-u-v has a perfect matching, and by conformality of $H$, the graph GuvG-u-v has a perfect matching. Consequently, by Proposition 1.2, at most one of $u$ and $v$ lies in $B$. ∎
@@
If there exist two or more components of GBG-B that contain branch vertices of $H$, then there exists a component of GBG-B that contains at least two branch vertices of $H$.
@@
Suppose to the contrary that at least two components of GBG-B contain branch vertices of $H$, but each such component contains exactly one branch vertex. Note that either $J=\theta$ or otherwise each vertex of $J$ has at least two neighbors. Thus, it follows from 2.9.1 that there exist components $K$ and $L$ of GBG-B that contain branch vertices $u$ and $v$, respectively, such that $u$ and $v$ are adjacent in $J$. This contradicts Lemma 2.3. ∎
We remark that if $J=\theta$, then the above statement implies that either both branch vertices belong to a particular component of GBG-B, or otherwise precisely one of them lies in $B$.
@@
There exists a component $L$ of GBG-B that contains at least |V(J)|1|V(J)|-1 branch vertices.
@@
The statement holds trivially for $J=\theta$. Otherwise, using 2.9.2, there exists a component $L$ of GBG-B that contains at least two branch vertices of $H$. Observe that $\partial(L)$ is a nontrivial tight cut; since $J$ is bicritical and cubic, we invoke Corollary 2.8 to infer that $\overline{V(L)}$ contains at most one branch vertex. ∎
@@
2.2 Handling 22-separations: a proof of Theorem 1.25
@@
Let $J$ and $G$ be matching covered graphs, and let $S$ denote a 22-separation of $G$. If some marked $S$-component of $G$ is $J$-based, then $G$ is also $J$-based.
@@
Let $S:=\{u,v\}$, and let $K$ and $L$ denote two distinct (even) components of GSG-S. Suppose the marked $S$-component $K^{\prime}$ (corresponding to $K$) is $J$-based, and let $H^{\prime}$ denote a conformal bisubdivision of $J$ in $K^{\prime}$. Let $e$ denote the marker edge of $K^{\prime}$, and let $P$ denote a conformal odd uvuv-path in $L$ provided by Proposition 1.22 (iii). Depending on whether or not $e$ belongs to $H^{\prime}$, either $H:=H^{\prime}+P-e$ or $H:=H^{\prime}$, respectively, is a conformal bisubdivision of $J$ in $G$. ∎
@@
Let $S:=\{u,v\}$ denote a 22-separation of a matching covered graph $G$, and let $J$ denote any 33-connected cubic graph. Assume that if $J=\theta$ then GSG-S has precisely two components. If $G$ is $J$-based then some marked $S$-component of $G$ is also $J$-based.
@@
We let $H$ denote a bisubdivision of $J$ that is conformal in $G$. Since $H$ is a subdivision of a 33-connected graph, and since $S$ is a 22-separation of $G$, we immediately conclude that there do not exist two components of GSG-S each of which contains a branch vertex of $H$. This proves the following.
@@
Since $\theta$ is the only 33-connected cubic graph of order two, if $J\neq\theta$ then there exists a unique marked $S$-component of $G$, denoted by $L^{\prime}$, that contains all branch vertices of $H$. On the other hand, if $J=\theta$ then, by our hypothesis, there are precisely two marked $S$-components; in this case, each of them may contain both branch vertices, and this happens if and only if $u$ and $v$ are the branch vertices of $H$. So, if $J=\theta$, we let $L^{\prime}$ denote the marked $S$-component that minimizes the number of $\partial_{G}(V(L^{\prime}))$-crossing paths. We let $e$ denote the marker edge of $L^{\prime}$, let $D$ denote the even cut $\partial_{G}(V(L^{\prime}))$, and let $M$ denote a perfect matching of GV(H)G-V(H).
@@
Suppose to the contrary that $H$ has at least two $D$-crossing paths, say $P$ and $Q$. Since all branch vertices belong to $V(L^{\prime})$, and since $S:=\{u,v\}$ is a 22-separation of $G$, each of $P$ and $Q$ contains $u$ as well as $v$. This immediately implies that $u$ and $v$ are branch vertices of $H$. Furthermore, since $J$ is 33-connected and cubic, we conclude that $J=\theta$. Consequently, $H=P\cup Q\cup R$, where $R$ is the third uvuv-path. Observe that if $K^{\prime}$ is the marked $S$-component of $G$ distinct from $L^{\prime}$, then the number of $\partial_{G}(V(K^{\prime}))$-crossing paths is at most one, and this contradicts our choice of $L^{\prime}$. ∎
@@
An edge $e$ of a matching covered graph $G$ is removable if GeG-e is also matching covered. In this case, if $G$ is a brick then GeG-e need not be a brick; in particular, GeG-e may have one or two vertices of degree two. In order to recover a brick, at the very least, we need to get rid of the degree two vertices (if any).
@@
Now, let $e$ be a removable edge of a brick $G$. As noted earlier, GeG-e has zero, one or two vertices of degree two. The graph $H$ obtained from GeG-e by bicontracting each of these degree two vertices is called the retract of GeG-e. Observe that $|V(G)|-|V(H)|\in\{0,2,4\}$; furthermore, each contraction vertex of $H$ (resulting from the bicontraction of degree two vertices), if any, has degree at least four. We say that $e$ is a thin edge of the brick $G$ if the retract of GeG-e is also a brick. For instance, the bicorn shown in Figure 8 has a unique removable edge that is also thin. On the other hand, if $e$ is any edge of the Petersen graph $\mathbb{P}$, the retract of $\mathbb{P}-e$ is precisely $T_{6}$; consequently, $\mathbb{P}$ is devoid of thin edges.
@@
In their work [4], Carvalho, Lucchesi and Murty defined four expansion operations — each of which comprises a sequence of zero, one or two bisplittings followed by adding an edge. In particular, if $e$ is a thin edge of a brick $G$, and $H$ is the retract of GeG-e, then $G$ may be obtained from $H$ using one of the four expansion operations. In our work, we do not require the precise definitions of these operations; the interested reader may refer to [4]. We shall simply use the facts that the constituent operations of adding an edge as well as of bisplitting preserve the $\theta$-based property; the former is easy to see, whereas the latter follows from Corollary 2.12. We are now ready to prove our characterization of $\theta$-free bricks restated below.
@@
By Proposition 1.12, $\mathbb{P}$ is $\theta$-free. Clearly, $K_{4}$ is $\theta$-free. Conversely, let $G$ be a $\theta$-free brick. We proceed by induction on the size of $G$. If $G\in\{K_{4},\overline{C_{6}},\mathbb{P}\}$, since $\overline{C_{6}}$ is $\theta$-based, the desired conclusion holds. Otherwise, by the Thin Edge Theorem (2.13), $G$ has a thin edge, say $e$, and let $H$ denote the retract of GeG-e. Clearly, GeG-e is $\theta$-free; by Corollary 2.12, $H$ is also $\theta$-free. Since $H$ is a brick, by the induction hypothesis, $H\in\{K_{4},\mathbb{P}\}$. Furthermore, $G$ is obtained from $H$ by one of the four expansion operations. Since $H$ is cubic, $G$ is obtained from $H$ by adding an edge. If $H$ is $K_{4}$, clearly $G$ is $\theta$-based; if $H$ is $\mathbb{P}$, by Proposition 1.12, $G$ is $\theta$-based; in either case, we have a contradiction. This proves Theorem 1.27. ∎